OpenStudy (vishal_kothari):

Are there positive irrational numbers a and b such that a b is rational..

OpenStudy (chriss):

\[\sqrt{2}*\sqrt{2} =2\] 2 is rational

OpenStudy (chriss):

Also, \[\sqrt{2}*\sqrt{8}=4\]

OpenStudy (vishal_kothari):

thanks...